#1. One lightyear is defined to be the distance light can travel in one year. (10 points)

What is this distance in meters?

How long does it take for light to get to the Moon?

#2. Multiply 1.783 x 102 * 4.4 x 10-3, taking into account significant figures. (10 points)

#3. What is the surface area of a sphere of diameter 2.4 x102 cm? (10 points)

#4. What is larger 4000 L or 4 x 105 cm3 (and you must show your reasoning.) (10 points)

#5. A car travels 120 meters in one direction in 20 seconds. Then the car returns ¾ of the way back in 10 seconds. (10 points)

Calculate the average speed of the car for the first part of the trip.

Find the average velocity of the car.

#6. A sports car moving at a constant speed travels 150 m in 4.00 s. If it then brakes and comes to a stop while decelerating at a rate of 6.0 m/s2 how long does it take to stop? (20 points)

#7. Explain a possible situation where you start with a positive velocity that decreases to a negative increasing velocity while there is a constant negative acceleration. (10 points)

#8. A stone is dropped from the roof of a high building. A second stone is dropped 1.25 s later. How long does it take for the stones to be 25.0 meters apart? (20 points)
Answer #1

(1)

time t = 1 year = 31536000 s

Distance d = v*t = 3*10^8 * 31536000

d = 9.4608*10^15 m

(2)

(1.783 * 10^2) * (4.4 * 10^(-3)) = 7.8452 * 10^(-1)

(3)

diameter = 2.4 *10^2 cm = 2.4 m

radius, r = d / 2 = 1.2 m

surface area of a sphere,

A = 4*pi*r^2 = 4*pi*(1.2)^2

A = 18.095 m^2

(4)

V1 = 4000 L

V2 = 4*10^5 cm^3 = 400 L (because 1 cm^3 = 0.001 L)

so, 4000 L is larger